OpenStudy (anonymous):

use basic identities to simplify: cosx-cosxsin^2x and cotxsecxsinx

geerky42 (geerky42):

cosx-cosxsin²x Factor cosx out. cosx(1-sin²x) = cosx(cos²x) = cos³x

geerky42 (geerky42):

cotxsecxsinx Knowing that \(\cot x = \dfrac{\cos x}{\sin x}\) and \(\sec x = \dfrac{1}{\cos x }\) cotxsecxsinx = 1

OpenStudy (anonymous):

thank you! what about proving this identity: (cscx-1)/(cscx+1) = (cot^2x)/(csc^2x+2cscx+1)

geerky42 (geerky42):

Factor csc²x+2cscx+1 Then multiply both sides by (cscx + 1)²